In a triangle side x is 3/5 if side z. If y is 3/4 of side x,. What is the ratio of y to z?


1
Expert's answer
2022-02-22T06:51:15-0500

Given, in a triangle side x is 3/5 of side z and y is 3/4 of side x.

This can be written in the equation as, "x= \dfrac{3}{5}z~\text{and } y = \dfrac{3}{4}x".

That is, "y = \dfrac{3}{4}\cdot\dfrac{3}{5}z=\dfrac{9}{20}z".

Therefore, "y:z=9:20."